find the missing coordinates 2x-3y=6:(3, ), ( ,-2), (-1, ), ( ,2)
Answer by umail08(87)

Find the missing coordinates 2x-3y=6:(3, ), ( ,-2), (-1, ), ( ,2)
2x-3y=6: (3,0), (0 ,-2), (-1, -8/3), (6,2)

If x=3, then:
2(3) – 3y = 6
6 – 3y = 6
-3y = 6 – 6
-3y = 0
-3y/-3 = 0/-3
y = 0

If x = -1, then:
2(-1) – 3y = 6
-2 – 3y = 6
-3y = 6 +2
-3y = 8
-3y/-3 = 8/-3
y = -8/3



If y=-2, then:
2x – 3(-2) = 6
2x + 6 = 6
2x = 6 - 6
2x = 0
2x/2 = 0/2
x = 0

If y=2, then:
2x – 3(2) = 6
2x - 6 = 6
2x = 6 + 6
2x = 12
2x/2 = 12/2
x = 6